OpenStudy (anonymous):

Solve the following system of equations by using the elimination method. x – 4y = 17 3x + 4y = 3

OpenStudy (anonymous):

Ok, add them together, the y will eliminate.

OpenStudy (anonymous):

times the first equation by 3 to get 3x.

OpenStudy (anonymous):

then solve for x.

OpenStudy (anonymous):

equation 1) x-4y=17 (times 3) 3x-12y=51 equation 2) 3x + 4y = 3 now solve for y by minusing 3x-3x=0 and when you get the x-value substitute it back into either equations to get the y value.